(-1, 2) (1, -1) What is the slope-intercept form?

(-1, 2) (1, -1)

What is the slope-intercept form?
Answer #1

please don't hesitate to give a "thumbs up" for the answer in case you're satisfied with it.

The slope intercept form is basically y = mx+c, where c is the y-intercept and m is the slope

Lets see.

Slope = y2-y1/x2-x1 = (-1-2)/1-(-1)) = -1.5

So, (y-2) = -1.5*(x-(-1))

y-2 = -1.5x -1.5

y = - 1.5x+0.5 <-- this is the correct answer.